Project

General

Profile

16/02/15

Attendance

Chris, Craig, Jeny, Michael, Paola, Nate, Nick, Dominick, Ryan

News (Matthew)

  • New ND geometry is available and there is a request for a sample to be simulated with it.
  • Issue with runNOvASAM.py - email thread about this - and a similar issue has hit calibration - looks to be resolved, but does it indicate the need for grid based testing?
  • Nate is moving on thanks for all your hard work over the years Nate. Ruth will step in for the short term and will help keep things moving while we identify a longer term replacement.

Ongoing simulation (Nate)

  • ND geant physics list: 4 done, only 1 left.
  • He is also working on documentation, could add to production procedures.

Ongoing reco (Satish)

  • FTS BIC jobs are done (2 errors)
  • Baseline files are done.
  • Next samples to start as soon as possible.
  • He is also in the processes of correcting some of our document

Ongoing PID (Bruno)

  • Bruno got approval for the production role from Gary, presumably this now has to be added by CD. He can start once this is all sorted out.

Reco keep-up (Jeny)

  • Friday ND keep-up is complete (only 2 files failed out of all them)
  • The FD was submitted last night, still running
  • Once this is complete she will work with Matthew to setup the cron jobs.

Calibration (Paola)

  • All failed due to the a problem with the release (see news). Matthew and Jonathan will help her resolve this.

Proposal for formalising production epochs (Matthew)

The consensus seemed to be that while it is good to formalise this process, doing it every two months might not be necessary or an optimal use of resources. That said, the March 1st date suggested does look to be a good one to start the next batch of files on. Some other points:

  • Chris requested that we name the current epochs: A1 A2 A3.

Discussion of FA branch - what is it good for? (all)

The general feeling that the FA branch isn't being used as we intended. Primarily because there was more development to do than was anticipated. As a results having two branches has led to some confusion and mistakes. Therefore we propose to retire this branch. Ahead of this Dominick will prepare a script that does the diff between FA and development to look for any changes that we should be aware of.

AOB

  • LEM production in FD cosmics is slow. Chris has talked to the Nu_e group and identified a loose preselection that buys us a factor of 4 - this needs to go into the PID. Dominick and Chris will work on putting this in place.